OpenStudy (anonymous):

For the line 4x-2y=6, what is the slope, y-intercept and x-intercept. Please show step by step.

OpenStudy (anonymous):

Simplifying 4x + -2y + -6 Reorder the terms: -6 + 4x + -2y Factor out the Greatest Common Factor (GCF), '2'. 2(-3 + 2x + -1y) Final result: 2(-3 + 2x + -1y) set equal to zero Simplifying -3 + 2x + -1y = 0 Solving -3 + 2x + -1y = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '3' to each side of the equation. -3 + 2x + 3 + -1y = 0 + 3 Reorder the terms: -3 + 3 + 2x + -1y = 0 + 3 Combine like terms: -3 + 3 = 0 0 + 2x + -1y = 0 + 3 2x + -1y = 0 + 3 Combine like terms: 0 + 3 = 3 2x + -1y = 3 Add 'y' to each side of the equation. 2x + -1y + y = 3 + y Combine like terms: -1y + y = 0 2x + 0 = 3 + y 2x = 3 + y Divide each side by '2'. x = 1.5 + 0.5y Simplifying x = 1.5 + 0.5y

OpenStudy (anonymous):

but slope cannot be determined because i think it is a parabola

OpenStudy (anonymous):

but you can try to plug in a random number and see what you get

OpenStudy (anonymous):

random number for x...to find y

OpenStudy (anonymous):

To find the y intercept, 4x-2y=6 4(0) -2y=6 -2y = 6 -2y/-2 = 6/-2 y = -3 to find the x intercept 4x-2y=6 4x-2(0)=6 4x = 6 4x/4 = 6/4 x = 6/4 or 1.5 Now use the two points to find the slope x intercept = (1.5,0) y intercept = (0,-3) m(slope) = y2-y1/x2-x1 m(slope) = -3-0/0-1.5 m(slope) = -3/-1.5 m(slope) = 2 Also to find the slope we can turn the equation into slop intercept form. y =mx+b This will also find the y intercept 4x-2y=6 -2y = -4x+6 -2y/-2 = -4x/-2 + 6/-2 y = 2x-3 slope is 2 and y intercept is -3 as shown in the equation